# Real $K$-Theory for $C^{*}$-Algebras: just the facts October 22, 2025

### v2

@@ 23. Examples and Calculations
As discussed in the previous section, we have (with $\{x\}$ designating the one-point space),
@@
$c$ 1 0 0 0 2 0 0 0
$r$ 2 0 1 0 1 0 0 0
$\eta$ 1 1 0 0 0 0 0 0
@@
$c$ $\left(\begin{smallmatrix}1\\ 1\end{smallmatrix}\right)$ 0 $\left(\begin{smallmatrix}1\\ 1\end{smallmatrix}\right)$ 0 $\left(\begin{smallmatrix}1\\ 1\end{smallmatrix}\right)$ 0 $\left(\begin{smallmatrix}1\\ 1\end{smallmatrix}\right)$ 0
$r$ $\left(\begin{smallmatrix}1&1\end{smallmatrix}\right)$ 0 $\left(\begin{smallmatrix}1&1\end{smallmatrix}\right)$ 0 $\left(\begin{smallmatrix}1&1\end{smallmatrix}\right)$ 0 $\left(\begin{smallmatrix}1&1\end{smallmatrix}\right)$ 0
$\eta$ 0 0 0 0 0 0 0 0
@@
With this notation, Examples 17.3 and 17.4 are generalized by the following:
@@
Furthermore, let $\tau^{p,q}$ be the involution on ${\bf{R}}^{n}$ (where $n=p+q$) given by the identity on $p$ coordinates and by multiplication by 1-1 on the remaining $q$ coordinates. Then
@@
We regard $S^{1}$ as the usual subset of the complex plane and let
@@
splits and we have an isomorphism
@@
The complex part $K_{*}(C^{{\mathbb{C}}}(S^{1}))$ is computed using the same split sequence. ∎
@@
is the corresponding real $C^{*}$-algebra.
@@
Since $1\in S^{1}$ is a fixed point of the involution $z\mapsto\overline{z}$, the condition that $\overline{f(z)}=f(\overline{z})$ implies that $f(1)\in{\mathbb{R}}$. So as in the previous calculation, we have a map $\mathrm{ev}_{1}\colon\real(iS^{1})\rightarrow{\mathbb{R}}$; this time the kernel is isomorphic to $V{\mathbb{R}}=\real_{0}(i{\mathbb{R}})$. Again, as in the previous calculation, the map splits yielding
@@
This follows from the Künneth Formula for real $K$-theory.
@@ 24. $KO$-theory with Coefficients
Let $G$ be a countable abelian group and let $N^{\mathbb{R}}_{G}$ denote any real $C^{*}$-algebra in the real bootstrap category such that
@@
For any countable abelian group $G$, there exists a real $C^{*}$-algebra $N_{G}^{\mathbb{R}}$ satisfying the given properties. Furthermore, the algebra $N_{G}^{\mathbb{R}}$ is unique in the sense that if $N_{G}^{\mathbb{R}}$ and $\widetilde{N}_{G}^{\mathbb{R}}$ both satisfy the given properties, then there is an isomorphism $KO_{*}(N_{G}^{\mathbb{R}})\cong KO_{*}(\widetilde{N}_{G}^{\mathbb{R}})$.
The stronger and more precise statement is that $N_{G}^{\mathbb{R}}$ and $\widetilde{N}_{G}^{\mathbb{R}}$ are $KK$-equivalent (via real $KK$-theory). That is, there is an element
@@
that implements the isomorphism stated in the Proposition 24.2. We will not prove this proposition since it requires the use of united $K$-theory which we have not developed. It relies on results from Section 8.4 of [42] and the Main Theorem of [17].
@@
$\lim_{n\to\infty}(N_{\mathbb{Z}}^{\mathbb{R}},\phi_{n})=N_{\mathbb{Q}}^{\mathbb{R}}\;.$
Then we also have
$\lim_{n\to\infty}(A\otimes N_{\mathbb{Z}}^{\mathbb{R}},\mathrm{id}\otimes\phi_{n})=A\otimes N_{\mathbb{Q}}^{\mathbb{R}}\;.$
@@
$=\lim_{n\to\infty}(KO_{*}(A\otimes N_{\mathbb{Z}}^{\mathbb{R}}),(\mathrm{id}\otimes\phi_{n})_{*})$
$=\lim_{n\to\infty}(KO_{*}(A),k_{n})$
@@
For the record, there are are three natural ways to obtain models for $N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}$.
@@
    Let $D^{2}$ be the unit disk and let $S^{1}$ be the unit circle in the complex plane. For $n\geq 2$, let $\zeta_{n}=\exp(2\pi j/n)\in S^{1}$ be the principal root of unity and define an equivalence relation on $S^{1}=\partial(D^{2})\subset{\mathbb{C}}$ generated by $z\sim\zeta_{n}\cdot z$. Then define the reduced lens space by
@@
    $0\rightarrow C_{0}(\rm{int}(D^{2}),{\mathbb{C}})\xrightarrow{\iota}C_{0}(L_{n},{\mathbb{C}})\xrightarrow{\rm{~ev~}}C_{0}(S^{1}-\{1\},{\mathbb{C}})\rightarrow 0$
@@
    $\partial_{1}\colon K_{1}(C(S^{1}-\{1\},{\mathbb{C}}))\rightarrow K_{0}(C_{0}(\rm{int}(D^{2}),{\mathbb{C}}))$
@@
    $0\rightarrow C_{0}(\rm{int}(D^{2}),{\mathbb{R}})\xrightarrow{\iota}C_{0}(L_{n},{\mathbb{R}})\xrightarrow{\rm{~ev~}}C_{0}(S^{1}-\{1\},{\mathbb{R}})\rightarrow 0$ (4)
@@
    $KO_{*}(C_{0}(S^{1}-\{1\},{\mathbb{R}}))=KO_{*}(S{\mathbb{R}})=\Sigma KO_{*}({\mathbb{R}})$
@@
    $KO_{*}(C_{0}(\rm{int}(D^{2})),{\mathbb{R}})=KO_{*}(S^{2}{\mathbb{R}})=\Sigma^{2}KO_{*}({\mathbb{R}})\ ;.$
@@
    $\partial_{1}\colon KO_{-1}(C(S^{1}-\{1\},{\mathbb{R}}))\rightarrow KO_{-2}(C_{0}(\rm{int}(D^{2})),{\mathbb{R}})$
@@
    $KO_{-2}(C_{0}(L_{n},{\mathbb{R}}))={\mathbb{Z}}_{n}\quad\text{and}\quad KO_{-3}((C_{0}(L_{n},{\mathbb{R}}))=0\;.$
    This is a double shift of what is specified in the definition, so to correct for this we take $N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}=V^{2}C_{0}(L_{n},{\mathbb{R}})$.
@@
    Then by the real $K$-theory calculations for $\mathcal{O}^{\mathbb{R}}_{n+1}$ in Theorem 1.6.8 of [76] and Section 5.1 of [15] we have $KO_{-1}(\mathcal{O}_{n+1}^{\mathbb{R}})=0$ and $KO_{0}(\mathcal{O}_{n+1}^{\mathbb{R}})={\mathbb{Z}}_{n}$. Therefore $N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}=\mathcal{O}^{\mathbb{R}}_{n+1}$. Note that by Theorem 10.2 of [20], any unital simple purely infinite model for $N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}$ is isomorphic to $\mathcal{O}_{n+1}^{\mathbb{R}}$ (this is stronger than what is stated in Proposition 24.2).
Theorem 24.7.
@@
such that the following Bockstein sequence is exact:
@@
The existence of $\rho_{n}$ follows from Theorem 24.5 based on the surjective homomorphism ${\mathbb{Z}}\rightarrow{\mathbb{Z}}_{n}$, but to obtain the exact sequence we take a different approach using the lens space model of $N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}$. Start with the short exact sequence shown in displayed Equation (4) and tensor with a real $C^{*}$-algebra $A$. Since commutative $C^{*}$-algebras are exact, the resulting sequence
@@
Then by using $C_{0}(L_{n},{\mathbb{R}})=S^{2}N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}$ this becomes
@@
This gives the natural transformations $\rho_{n}=i_{*}$ and $\beta_{n}=\mathrm{ev}_{*}$. The differential $\partial\colon KO_{*}(A)\rightarrow KO_{*}(A)$ is known from [75] to be multiplication by $n$ in the case of $K_{0}({\mathbb{C}})={\mathbb{Z}}$ for complex $K$-theory. Then it must be multiplication by $n$ in the real case $KO_{0}({\mathbb{R}})$ since there is an isomorphism
@@
It follows that $\partial$ is multiplication by $n$ for any $A$ in the bootstrap class, using the Universal Coefficient Theorem for real $C^{*}$-algebras in [16]. ∎
@@
For the long exact sequence, we use the lens space model of $N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}$. There is an obvious injection $L_{nm}\rightarrow L_{n}$ which induces a homomorphism
$\alpha\colon N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}\rightarrow N_{{\mathbb{Z}}_{mn}}^{\mathbb{R}}\;.$
Now we apply the mapping cone construction (Definition 4.9) to get the short exact sequence
@@
Note that the partial map arising from this long exact sequence coincides with the map $\alpha$ by Theorem 3.5 of [74], which applies in the case of real $K$-theory just as for complex $K$-theory. Using the fact that $\alpha_{*}$ is multiplication by $m$ on
$KO_{0}(N_{{\mathbb{Z}}_{n}}^{\mathbb{R}})\rightarrow KO_{0}(N_{{\mathbb{Z}}_{mn}}^{\mathbb{R}})\quad\text{(which is ${\mathbb{Z}}_{n}\rightarrow{\mathbb{Z}}_{mn}$)}$
@@
$KO_{i}(N_{{\mathbb{Z}}_{mn}}^{\mathbb{R}})\rightarrow KO_{i+1}(C\alpha)$
@@
$0\rightarrow K^{\mathcal{CRT}}(A)\otimes_{\mathcal{CRT}}\mathcal{M}_{G}\xrightarrow{\mu}K^{\mathcal{CRT}}(A;G)\xrightarrow{\partial}\rm{Tor}_{\mathcal{CRT}}(K^{\mathcal{CRT}}(A),\mathcal{M}_{G})\rightarrow 0\;$
@@
25. $K$-theory and Unitaries

### v3

@@ 23. Examples and Calculations
Let $\{x\}$ designate the one-point space. As discussed in the previous section, we have
@@
$c_{n}$ 1 0 0 0 2 0 0 0
$r_{n}$ 2 0 1 0 1 0 0 0
$\eta_{n}$ 1 1 0 0 0 0 0 0
This table shows the groups of real and complex $K$-theory, as well as the natural transformations
$c_{n}\colon$ $KO_{n}({\mathbb{R}})\rightarrow K_{n}({\mathbb{R}}_{\mathbb{C}})$
$r_{n}\colon$ $K_{n}({\mathbb{R}}_{\mathbb{C}})\rightarrow KO_{n}({\mathbb{R}})$
$\eta_{n}\colon$ $KO_{n}({\mathbb{R}})\rightarrow KO_{n+1}({\mathbb{R}})\;.$
In the case at hand, each group is generated by a single element, so we can express the homomorphism by a single integer. For example, the table shows that
$c_{4}\colon KO_{4}({\mathbb{R}})\rightarrow KO_{4}({\mathbb{R}}_{C})$
is the map from $\mathbb{Z}$} to $\mathbb{Z}$} given by multiplication by 2. In the examples below we use matrices to represent the homomorphisms when one of the groups has more than one generator.
@@
$c_{n}$ $\left(\begin{smallmatrix}1\\ 1\end{smallmatrix}\right)$ 0 $\left(\begin{smallmatrix}1\\ -1\end{smallmatrix}\right)$ 0 $\left(\begin{smallmatrix}1\\ 1\end{smallmatrix}\right)$ 0 $\left(\begin{smallmatrix}1\\ -1\end{smallmatrix}\right)$ 0
$r_{n}$ $\left(\begin{smallmatrix}1&1\end{smallmatrix}\right)$ 0 $\left(\begin{smallmatrix}1&-1\end{smallmatrix}\right)$ 0 $\left(\begin{smallmatrix}1&1\end{smallmatrix}\right)$ 0 $\left(\begin{smallmatrix}1&-1\end{smallmatrix}\right)$ 0
$\eta_{n}$ 0 0 0 0 0 0 0 0
@@
With this notation, Examples 17.3 and 17.4 may be generalized as follows:
@@
Now let $n=p+q$ and let $\tau^{p,q}$ be the involution on ${\bf{R}}^{n}$ given by the identity on $p$ coordinates and by multiplication by 1-1 on the remaining $q$ coordinates. Then
@@
Regard $S^{1}$ as the usual subset of the complex plane and let
@@
splits, and we have an isomorphism
@@
The complex groups $K_{*}(C^{{\mathbb{C}}}(S^{1}))$ are computed using the same split sequence. ∎
@@
is the corresponding real $C^{*}$-algebra. Then we have:
@@
Since $1\in S^{1}$ is a fixed point of the involution $z\mapsto\overline{z}$, the condition that $\overline{f(z)}=f(\overline{z})$ implies that $f(1)\in{\mathbb{R}}$. So as in the previous calculation, we have a map $\mathrm{ev}_{1}\colon\real(iS^{1})\rightarrow{\mathbb{R}}$; this time the kernel is isomorphic to $V{\mathbb{R}}=\real_{0}(i{\mathbb{R}})$. So there is a split short exact sequence
$0\rightarrow\real_{0}(i{\bf{R}})\rightarrow\real(iS^{1})\xrightarrow{\mathrm{ev}_{1}}{\mathbb{R}}\rightarrow 0\;.$
Again, as in the previous calculation, the fact that the sequence splits implies that
@@
This follows as in Example 23.7 and  23.8 using the split short exact sequences
$0\rightarrow\real_{0}({\bf{R}}^{n})\rightarrow\real(S^{n})\xrightarrow{\mathrm{ev}_{1}}{\mathbb{R}}\rightarrow 0\;$
and
$0\rightarrow\real_{0}(i{\bf{R}}^{n})\rightarrow\real(iS^{n})\xrightarrow{\mathrm{ev}_{1}}{\mathbb{R}}\rightarrow 0\;.$
@@ 24. $KO$-theory with Coefficients
Let $G$ be a countable abelian group and let $N^{\mathbb{R}}_{G}$ denote any real $C^{*}$-algebra in the pre-bootstrap category such that
@@
Let $G$ be a countable abelian group. Then
  1. (1)
    There exists a real $C^{*}$-algebra $N^{\mathbb{R}}_{G}$ in the pre-bootstrap category satisfying the three conditions listed above.
  2. (2)
    The real $C^{*}$-algebra $N^{\mathbb{R}}_{G}$ is unique in the sense that if there are two real $C^{*}$ algebras $N^{\mathbb{R}}_{G}$ and $\tilde{N}^{\mathbb{R}}_{G}$, each satisfying the three conditions listed above, then there is an isomorphism
    $KO_{*}(N^{\mathbb{R}}_{G})\cong KO_{*}(\tilde{N}^{\mathbb{R}}_{G}).$
  3. (3)
    In fact, $N_{G}^{\mathbb{R}}$ and $\widetilde{N}_{G}^{\mathbb{R}}$ are $KK$-equivalent (via real $KK$-theory). That is, there is an element
@@
    that implements the isomorphism stated in (2).
We will not prove this proposition since it requires the use of united $K$-theory which we have not developed. It relies on results from Section 8.4 of [42] and the Main Theorem of [17].
@@
[…]
up to $KK$-equivalence. Then we also have
[…]
@@
[…]
@@
Corollary 24.7.
Suppose that $G$ is a countable abelian torsion-free 2-divisible group. Then $K$-theory with coefficients in $G$ has period 4. In particular, there are isomorphisms
$KO_{i}(A;{\mathbb{Q}})=KO_{i+4}(A;{\mathbb{Q}})\qquad\text{and}\qquad KO_{i}(A;{\mathbb{Z}}[\tfrac{1}{2}])=KO_{i+4}(A;{\mathbb{Z}}[\tfrac{1}{2}])\;.$
Proof.
The proof of this Corollary requires the full Künneth Formula and results about $\mathcal{CRT}$-modules. Here we will only prove the Corollary in the case $G={\mathbb{Q}}$ using Theorem 24.6 above. In fact, we will prove that there is an isomorphism
$KO_{i}(A)\otimes G\cong KO_{i+4}(A)\otimes G$
where $G$ is torsion-free and 2-divisible, which implies the result for $G={\mathbb{Q}}$ using the isomorphism $KO_{i}(A;{\mathbb{Q}})\cong KO_{i}(A)\otimes{\mathbb{Q}}$.
Consider the natural transformation
$\xi\colon KO_{i}(A)\rightarrow KO_{i+4}(A)$
given by multiplication by $\xi\in KO_{4}({\mathbb{R}})$. We know that $\xi^{2}=4\beta_{O}$, by Theorem 22.2, where $\beta_{O}$ is the real Bott periodicity isomorphism. This implies that
$(\xi\otimes\mathrm{id})^{2}\colon KO_{i}(A)\otimes G\rightarrow KO_{i+8}(A)\otimes G$
is given by multiplication by 4. Then the hypothesis on $G$ imply that $(\xi\otimes\mathrm{id})^{2}$ is an isomorphism, and this implies that $\xi\otimes\mathrm{id}$ is an isomorphism. ∎
@@
There are are three natural ways to obtain models for $N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}$. Here they are.
@@
    Let $D^{2}$ be the unit disk and let $S^{1}$ be the unit circle in the complex plane. For $n\geq 2$, let $\zeta_{n}=\exp(2\pi j/n)\in S^{1}$ be the principal $n$th root of unity and define an equivalence relation on $S^{1}=\partial(D^{2})\subset{\mathbb{C}}$ generated by $z\sim\zeta_{n}\cdot z$. Then define the reduced lens space by
@@
    $0\rightarrow C^{{\mathbb{C}}}_{0}(\rm{int}(D^{2}))\xrightarrow{\iota}C^{{\mathbb{C}}}_{0}(L_{n})\xrightarrow{\rm{~ev~}}C^{{\mathbb{C}}}_{0}(S^{1}-\{1\})\rightarrow 0$
@@
    $\partial_{1}\colon K_{1}(C^{{\mathbb{C}}}_{0}(S^{1}-\{1\}))\rightarrow K_{0}(C^{{\mathbb{C}}}_{0}(\rm{int}(D^{2}))$
@@
    $0\rightarrow\real_{0}(\rm{int}(D^{2}))\xrightarrow{\iota}\real_{0}(L_{n})\xrightarrow{\rm{~ev~}}\real_{0}(S^{1}-\{1\})\rightarrow 0$ (24.1)
@@
    $KO_{*}(\real_{0}(S^{1}-\{1\}))=KO_{*}(S{\mathbb{R}})=\Sigma KO_{*}({\mathbb{R}})$
@@
    $KO_{*}(\real_{0}(\rm{int}(D^{2})))=KO_{*}(S^{2}{\mathbb{R}})=\Sigma^{2}KO_{*}({\mathbb{R}})\ ;.$
@@
    $\partial_{1}\colon KO_{-1}(\real_{0}(S^{1}-\{1\}))\rightarrow KO_{-2}(\real_{0}(\rm{int}(D^{2})))$
@@
    $KO_{-2}(\real_{0}(L_{n}))={\mathbb{Z}}_{n}\quad\text{and}\quad KO_{-3}((\real_{0}(L_{n}))=0\;.$
    This is a double shift of what is specified in the definition, so to correct for this we take $N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}=V^{2}\real_{0}(L_{n})$.
@@
    Then by the real $K$-theory calculations for $\mathcal{O}^{\mathbb{R}}_{n+1}$ in Theorem 1.6.8 of [76] and Section 5.1 of [15] we have $N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}=\mathcal{O}^{\mathbb{R}}_{n+1}$. Note that by Theorem 10.2 of [20], any unital simple purely infinite model for $N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}$ is isomorphic to $\mathcal{O}_{n+1}^{\mathbb{R}}$ (this is stronger than what is stated in Proposition 24.2).
The next two results generalize Proposition 1.6 in [75] (the Bockstein exact sequence) and Part (2) of Proposition 2.6 in [75] from the complex case to the real case.
Proposition 24.8.
@@
such that the following Bockstein sequence is exact
@@
The existence of $\rho_{n}$ follows from Theorem 24.5 based on the surjective homomorphism ${\mathbb{Z}}\rightarrow{\mathbb{Z}}_{n}$, but to obtain $\beta_{n}$ and the exact sequence we take a different approach using the lens space model of $N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}$. Start with the short exact sequence shown in displayed Equation (24.1) above and tensor with a real $C^{*}$-algebra $A$. Since commutative $C^{*}$-algebras are exact, the resulting sequence
@@
Now we have $N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}=S^{-2}\real_{0}(L_{n})$, so
$KO_{i}(A;{\mathbb{Z}}_{n})=KO_{i}(A\otimes N_{{\mathbb{Z}}_{n}}^{\mathbb{R}})=KO_{i-2}(C_{0}(L_{n},A))\;.$
Using this our long exact sequence can be written as
@@
This gives the natural transformations $\rho_{n}=i_{*}$ and $\beta_{n}=\mathrm{ev}_{*}$. It remains to show that the differential $\partial\colon KO_{*}(A)\rightarrow KO_{*}(A)$ is multiplication by $n$. The map arising from the same construction in the case of complex $K$-theory is known to be multiplication by $n$ in for complex $K$-theory (see Lemma 1.4 of [75]). In particular, consider the $\partial$ maps when $A={\mathbb{R}}$ and when $A={\mathbb{C}}$ which commute with the complexification map as in the diagram below, where each group is isomorphic to $\mathbb{Z}$} and the vertical maps are isomorphisms. Since the map $\partial$ is multiplication by $n$ on $K_{0}({\mathbb{C}})$, the same is true for $\partial$ on $KO_{0}({\mathbb{R}})$.
${KO_{0}({\mathbb{R}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\partial}$$\scriptstyle{c_{0}}$${KO_{0}({\mathbb{R}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{c_{0}}$${K_{0}({\mathbb{C}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\partial}$${K_{0}({\mathbb{C}})}$
Then it must also be multiplication by $n$ in the real case $KO_{0}({\mathbb{R}})$ since there is an isomorphism
@@
Now for any real $C^{*}$-algebra $A$ in the pre-bootstrap category, we have
$KO_{i}(A\otimes{\mathbb{R}})\cong KO_{i}(A)\otimes KO_{0}({\mathbb{R}})$
by a special case of the Künneth Formula in [15]; and furthermore by naturality of the Künneth Formula, the diagram below commutes
${KO_{i}(A\otimes{\mathbb{R}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\partial}$${KO_{i}(A)\otimes KO_{0}({\mathbb{R}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{1\otimes\partial}$${KO_{i}(A\otimes{\mathbb{R}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${KO_{i}(A)\otimes KO_{0}({\mathbb{R}})}$
where the horizontal maps are isomorphisms of the Künneth Theorem. Since the vertical map $1\otimes\partial$ is multiplication by $n$, it then follows that $\partial$ is multiplication by $n$ for any $A$ in the pre-bootstrap category. ∎
@@
For the long exact sequence, we use the lens space model of $N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}$. There is an obvious quotient map $L_{n}\rightarrow L_{mn}$ which induces a injective homomorphism
$\alpha\colon N_{{\mathbb{Z}}_{mn}}^{\mathbb{R}}\rightarrow N_{{\mathbb{Z}}_{n}}^{\mathbb{R}}\;.$
It can be show that the map induced by $\alpha$ on real $K$-theory
$KO_{0}(N_{{\mathbb{Z}}_{mn}}^{\mathbb{R}})\rightarrow KO_{0}(N_{{\mathbb{Z}}_{n}}^{\mathbb{R}})\;,$
is the natural surjective map from ${\mathbb{Z}}_{mn}$ to ${\mathbb{Z}}_{n}$. (This is also true of the map on complex $K$-theory.) Now we apply the mapping cone construction (Definition 4.9) to get the short exact sequence
@@
Note that the partial map arising from this long exact sequence coincides with the map $\alpha_{*}$ by Theorem 3.5 of [74], which applies in the case of real $K$-theory just as for complex $K$-theory. Using the fact that $\alpha_{*}$ is surjective on
$KO_{0}(N_{{\mathbb{Z}}_{mn}}^{\mathbb{R}})\rightarrow KO_{0}(N_{{\mathbb{Z}}_{n}}^{\mathbb{R}})\quad\text{(which is ${\mathbb{Z}}_{mn}\rightarrow{\mathbb{Z}}_{n}$)}$
@@
$KO_{i}(N_{{\mathbb{Z}}_{mn}}^{\mathbb{R}})\rightarrow KO_{i-1}(C\alpha)$
@@
$0\rightarrow K^{\mathcal{CRT}}(A)\otimes_{\mathcal{CRT}}K^{\mathcal{CRT}}(N_{G})\xrightarrow{\mu}K^{\mathcal{CRT}}(A;G)\xrightarrow{\partial}\rm{Tor}_{\mathcal{CRT}}(K^{\mathcal{CRT}}(A),K^{\mathcal{CRT}}(N_{G}))\rightarrow 0\;$
@@
25. $K$-theory, KOKO-theory, and Unitaries